graph the Linear Equation 16x-4y=28
Answer by Alan3354(69443)

graph the Linear Equation 16x-4y=28
---------
16x-4y=28
Solve for y for convenience
4x - y = 7
y = 4x - 7
Pick 2 values for x and find y.
Plot the 2 points.
Draw a line thru them.